# All-sky search for individual Primordial Black Hole bursts with LHAASO Zhen Cao

### v2

@@
All-sky search for individual Primordial Black Hole bursts with LHAASO
@@ Abstract
A PBH with an initial mass of approximately g is expected to explode today in a final burst of Hawking radiation. In this work, we conduct an all-sky search for individual PBH burst events using the data collected from March 2021 to July 2024 by the Water Cherenkov Detector Array of the Large High Altitude Air Shower Observatory (LHAASO). Three PBH burst durations, 10 s, 20 s, and 100 s, are searched, with no significant PBH bursts observed. The upper limit on the local PBH burst rate density is set to be as low as 181 pc-3 yr-1 at 99 confidence level, representing the most stringent limit achieved to date.
Primordial black holes (PBHs) are predicted to produce from overdensities during the early universe through various mechanisms [1, 2, 3]. The PBH mass is related to the cosmological horizon mass at formation, spanning an enormous range from the Planck mass ( g) to solar mass or larger. PBHs have attracted much attention recently, as they could be responsible for some of the gravitational wave signals from binary black holes detected by the LIGO/Virgo [4, 5, 6].
As PBHs formed in the radiation-dominated era, they are non-baryonic and could serve as a good dark matter candidate. Over the past decades, the abundance of PBHs has been tightly constrained by numerous cosmological and astrophysical observations.
@@
As an important probe of the early Universe, the observational evidence of a PBH final evaporation event could give significant information about the early Universe, high energy particle physics beyond the energy scale of the Large Hadron Collider [17], and the convolution of gravitation with thermodynamics. In particular, the final burst should also produce new particles predicted in several extensions of the SM of particle physics. The existence of new particles could significantly modify the final evaporation, making its detection a unique window to explore physics beyond the SM [18, 20, 19]. Even the non-existence of PBHs also yields important constraints about the early Universe [12]. For instance, limits on the local burst rate density of PBHs can provide tighter limits on the power spectrum of primordial density perturbations at much smaller scales than those probed by the cosmic microwave background [21], thereby allowing the exploration of various inflation models.
@@
Any burst detected in the data will be further validated as a PBH candidate by comparing the observed spectrum and light curves with the expected theoretical ones for PBHs to verify consistency.
@@
[34] is a multipurpose and comprehensive extensive air shower array, located at Mt. Haizi at an altitude of 4410 meters above sea level in Daocheng, Sichuan province, China. It is designed for the study of cosmic-rays and gamma-rays with energies from sub-TeV to beyond 1 PeV. LHAASO is composed of three sub-arrays: the KiloMeter Squared Array (KM2A) for gamma-ray astronomy above 10 TeV and cosmic ray physics; WCDA for gamma-rays above a few hundred GeV, and the Wide Field-of-view air Cherenkov Telescopes Array (WFCTA) for cosmic ray physics from 10 TeV to 1 EeV.
@@
The expected number of gamma-rays () detected by the WCDA as a function of the PBH distance , , and zenith angle is given by:
where and are the WCDA’s energy range, denotes the effective area of WCDA as a function of photon energy and zenith angle. is simulated with the G4WCDA, a geant4-based [38] package developed to simulate the WCDA response.
Assuming a uniform distribution of PBHs in the solar neighborhood, we simulate a random set of PBH burst events under the assumption of a local burst rate density , with distances up to . No observable signals are expected beyond , even for the largest burst duration of . In the WCDA energy range, the burst emission time profile of PBHs is calculated to be (according to the method described in Section 3.5 of Ref. [15]), which is used to simulate the light curve of each PBH burst. As the remaining lifetime decreases, the emission rate increases sharply during the final few seconds, resulting in a VHE gamma-ray burst. The spatial distribution of each burst event is simulated by convolving the expected point signal with the angular resolution of the detector, also known as the point spread function (PSF).
@@
For each shower event, the root mean square distance (RMDS) for shower core fitting is calculated based on the 10 hottest detectors. To achieve better resolution of the shower core position, the RMDS is required to be m. The selected events are subsequently categorized into seven groups based on the effective number of triggered PMT units, denoted as , with ranges: [30–60), [60–100), [100–200), [200–300), [300–500), [500–800), and 800. To suppress the cosmic-ray background, a gamma-proton discrimination parameter, [39], is utilized to efficiently select gamma-like events.
@@
The values are required to be less than 1.02, 0.90, 0.88, 0.88, 0.84, 0.84, and 0.84 for seven groups, respectively. These criteria are optimized based on the observed significance of the Crab Nebula. Depending on the energy of the gamma rays and cosmic rays, these criteria can retain roughly more than 60 of gamma-rays while suppressing over 97 of the cosmic-ray background. Following this background suppression, approximately 44 million gamma-like events were observed by WCDA each day.
@@
The PBH burst search is conducted by scanning all gamma-like events across three search timescales ( s, 20 s, and 100 s) for significant excesses over the expected cosmic-ray background. All the surviving gamma-like events in each search time duration are filled into the sky map in celestial coordinates (R.A. and decl.), with the sky map divided into pixels, according to their arrival direction. The spatial scan for a sky map is performed by sliding a rectangular grid across the sky map until the entire WCDA field of view is covered. The grid size is optimized to based on the sensitivity study of PBH bursts, using the PBH signal MC sample and a background MC sample generated through Poisson sampling from the estimated background map. The width of the spatial grid in R.A. is scaled with decl. according to to account for the smaller sky sphere as the decl. moves away from . We continuously slide the time window and repeat the spatial scan until the entire WCDA dataset has been scanned. The step size for sliding the spatial grid (or temporal window) is set to be smaller than the grid size (or the duration of the time window) to ensure a high probability of capturing the majority of gamma-ray events from an individual PBH burst. The total number of search grids ranges from to , depending on the length of the search timescales.
@@
The cosmic-ray background at each spatial grid is estimated using the ‘direct integral method’ [40]. This method assumes that the detection efficiency varies slowly and the time average within a relatively short time window can properly reflect the background. In this analysis, we use a 3-hour integration time to estimate the detector acceptance. To minimize statistical fluctuations in background estimation, we require at least 10 events in the corresponding acceptance histogram position for each search grid. However, this requirement is often not met for large groups due to the low instantaneous event rate, and thus reduces the significance of the PBH burst signal. Therefore, we combine the hit groups with and re-categorize the data into three groups with in ranges of and .
@@
We performed a likelihood-ratio test to assess the significance of a burst event across three groups within the same grid location. The total test statistics (TS) for each search grid location is calculated as follows:
@@
All grids in the data scan with a significance greater than (corresponding to the -value ) are recorded for further identification of the PBH burst signals. The -value distribution for the recorded grids is expected to consist of two components: cosmic-ray background fluctuations and potential PBH bursts in the data.
@@
GRB 221009A is an extraordinarily bright burst and has been observed by LHAASO [41, 42]. In this work, we also detect GRB 221009A significantly using the blind scan method (details can be found in the Supplemental Material [33]).
@@
As no significant PBH burst signal is observed in the -value distribution of the data, the upper limit on the local PBH burst rate density is determined at the 99 confidence level (CL). In our search, the nearby grids have approximately 80 overlap in both spatial and temporal searches. This overlap causes the grids to be non-independent, altering the statistical behavior of each -value bin in Fig. 1.
@@
Hence, we select three low-background-level -value regions as a single -value bin for the three search durations to determine the upper limits of the local PBH burst density. The chosen -value regions are for s, for s, and for s, respectively. These ranges are optimized for the best sensitivity of PBH signals based on expected background counts and the simulated signal PDF. The details are described in the Supplemental Material [33].
The probability of the observed grid counts in the three chosen -value regions is calculated by summing over the probabilities of all possible signal and background count combinations. Therefore, the likelihood with a local burst rate density is given by:
where is the observed count of grids in the given -value regions for data. represents the simulated signal probability for signal grid counts under the local burst rate density of , and denotes the probability for the remaining background counts.
To determine the 99 CL upper limits of the local burst rate density, we perform a likelihood scan under various hypotheses for the local burst rate density. After incorporating systematic uncertainty, the 99 CL upper limits on the local PBH burst rate density are listed in Tab. 1 and shown in Fig. 2. Compared to previous constraints from other telescopes and observatories, LHAASO demonstrates the highest sensitivity to timescales ranging from several seconds to several hundred seconds.
@@ 99 CL upper limit (pc-3 yr-1)
For background estimation, the background in each grid is estimated using a 3-hour integration time. To assess the uncertainty in the background histograms, the integration time is varied from 3 to 5 hours. Furthermore, we use the -value range of as the fiducial region to verify the consistency between the data and the background simulation. The systematic uncertainty from the background estimation is estimated to be 5 across the three burst timescales.
In the PBH burst simulation, the atmospheric model in the MC simulation, along with variations in atmospheric conditions during operation, affects the detection efficiency, which is not properly modeled in the simulation. This uncertainty is estimated to be as large as 8 on the effective area [43]. Additionally, uncertainties in the PSF influence the number of signals within the given grids, thus impacting the -value distribution of PBH signals. The uncertainty in the simulated PSF is estimated to be by comparing the observed event profiles among Crab, Mrk 421, and Mrk 501 [43]. To estimate the systematic uncertainty due to the effective area and PSF simulation, we re-simulate the PBH burst signals by scaling the effective area by 92 and enlarging the PSF by . The total systematic uncertainty is accounted for by varying the background level within its corresponding uncertainty and using the re-simulated signal PDFs to repeat the upper limit measurement. The larger upper limit, compared to the nominal one, is adopted as the final result.

### v3

# All-Aky Search for Individual Primordial Black Hole Bursts with LHAASO Zhen Cao

@@
All-Aky Search for Individual Primordial Black Hole Bursts with LHAASO
@@ Abstract
A PBH with an initial mass of approximately g is expected to explode today in a final burst of Hawking radiation. In this work, we conduct an all-sky search for individual PBH burst events using the data collected from March 2021 to July 2024 by the Water Cherenkov Detector Array of the Large High Altitude Air Shower Observatory (LHAASO). Three PBH burst durations, 10 s, 20 s, and 100 s, are searched in the gamma-ray energy range from about 1 TeV to 15 TeV, with no significant PBH bursts observed. The upper limit on the local PBH burst rate density is set to be as low as 181 pc-3 yr-1 at 99 confidence level, representing the most stringent limit achieved to date.
Primordial black holes (PBHs) are predicted to form from overdensities during the early Universe through various mechanisms [1, 2, 3]. The PBH mass is related to the cosmological horizon mass at formation, spanning an enormous range from the Planck mass ( g) to solar mass or larger. PBHs have attracted much attention recently, as they could be responsible for some of the gravitational wave signals from binary black holes detected by the LIGO/Virgo [4, 5, 6].
As PBHs formed in the radiation-dominated era, they are nonbaryonic and could serve as a good dark matter candidate. Over the past decades, the abundance of PBHs has been tightly constrained by numerous cosmological and astrophysical observations.
@@
As an important probe of the early Universe, the observational evidence of a PBH final evaporation event could give significant information about the early Universe, high energy particle physics beyond the energy scale of the Large Hadron Collider [17], and the convolution of gravitation with thermodynamics. In particular, the final burst should also produce new particles predicted in several extensions of the SM of particle physics. The existence of new particles could significantly modify the final evaporation, making its detection a unique window to explore physics beyond the SM [18, 20, 19]. Even the nonexistence of PBHs also yields important constraints about the early Universe [12]. For instance, limits on the local burst rate density of PBHs can provide tighter limits on the power spectrum of primordial density perturbations at much smaller scales than those probed by the cosmic microwave background [21], thereby allowing the exploration of various inflation models.
@@
Any burst detected in the data is further validated as a PBH candidate by comparing the observed spectrum and light curves with the expected theoretical ones for PBHs to verify consistency.
@@
[34] is a multipurpose and comprehensive extensive air shower array, located at Mt. Haizi at an altitude of 4410 meters above sea level in Daocheng, Sichuan province, China. It is designed for the study of cosmic rays and gamma rays with energies from sub-TeV to beyond 1 PeV. LHAASO is composed of three subarrays: the KiloMeter Squared Array (KM2A) for gamma-ray astronomy above 10 TeV and cosmic ray physics; WCDA for gamma-rays above a few hundred GeV, and the Wide Field-of-View air Cherenkov Telescopes Array for cosmic ray physics from 10 TeV to 1 EeV.
@@
The expected number of gamma rays () detected by the WCDA as a function of the PBH distance , , and zenith angle is given by:
where and are the WCDA’s energy range, and denotes the effective area of WCDA as a function of photon energy and zenith angle. is simulated with the G4WCDA, a geant4-based [38] package developed to simulate the WCDA response.
Assuming a uniform distribution of PBHs in the solar neighborhood, we simulate a random set of PBH burst events under the assumption of a local burst rate density , with distances up to . No observable signals are expected beyond , even for the largest burst duration of . In the WCDA energy range, the burst emission time profile of PBHs is calculated to be (according to the method described in Sec. 3.5 of Ref. [15]), which is used to simulate the light curve of each PBH burst. As the remaining lifetime decreases, the emission rate increases sharply during the final few seconds, resulting in a VHE gamma-ray burst. The spatial distribution of each burst event is simulated by convolving the expected point signal with the angular resolution of the detector, also known as the point spread function (PSF).
@@
For each shower event, the root mean square distance (RMSD) for shower core fitting is calculated based on the 10 hottest detectors. To achieve better resolution of the shower core position, the RMSD is required to be m. The selected events are subsequently categorized into seven groups based on the effective number of triggered PMT units, denoted as , with ranges: [30 60), [60 100), [100 200), [200 300), [300 500), [500 800), and 800. The mean energies corresponding to the groups are approximately 1.0, 1.5, 2.1, 2.7, 3.8, 7.1, and 15.0 TeV, simulated using the energy spectrum of s.
To suppress the cosmic-ray background, a gamma-proton discrimination parameter, [39], is utilized to efficiently select gamma-like events.
@@
The values are required to be less than 1.02, 0.90, 0.88, 0.88, 0.84, 0.84, and 0.84 for seven groups, respectively. These criteria are optimized based on the observed significance of the Crab Nebula. Depending on the energy of the gamma rays and cosmic rays, these criteria can retain roughly more than 60 of gamma rays while suppressing over 97 of the cosmic-ray background. Following this background suppression, approximately 44 million gamma-like events are observed by WCDA each day.
@@
The PBH burst search is conducted by scanning all gamma-like events across three search timescales ( s, 20 s, and 100 s) for significant excesses over the expected cosmic-ray background. All the surviving gamma-like events in each search time duration are filled into the sky map in celestial coordinates (R.A. and decl.), with the sky map divided into square pixels, according to their arrival direction. The spatial scan for a sky map is performed by sliding a rectangular grid across the sky map until the entire WCDA field of view is covered. The grid size is optimized to based on the sensitivity study of PBH bursts using the PBH signal MC sample and a background MC sample generated through Poisson sampling from the estimated background map. The width of the spatial grid in R.A. is scaled with decl. according to to account for the smaller sky sphere as the decl. moves away from . We continuously slide the time window and repeat the spatial scan until the entire WCDA dataset has been scanned. The step size for sliding the spatial grid (or temporal window) is set to be smaller than the grid size (or the duration of the time window) to ensure a high probability of capturing the majority of gamma-ray events from an individual PBH burst. The total number of search grids ranges from to , depending on the length of the search timescales.
@@
The cosmic-ray background at each spatial grid is estimated using the ‘direct integral method’ [40]. This method assumes that the detection efficiency varies slowly and the time average within a relatively short time window can properly reflect the background. In this analysis, we use a three-hour integration time to estimate the detector acceptance. To minimize statistical fluctuations in background estimation, we require at least 10 events in the corresponding acceptance histogram position for each search grid. However, this requirement is often not met for large groups due to the low instantaneous event rate, and thus, it reduces the significance of the PBH burst signal. Therefore, we combine the hit groups with and recategorize the data into three groups with in ranges of and .
@@
We perform a likelihood-ratio test to assess the significance of a burst event across three groups within the same grid location. The total test statistics (TS) for each search grid location is calculated as follows:
@@
All grids in the data that scan with a significance greater than (corresponding to the -value ) are recorded for further identification of the PBH burst signals. The -value distribution for the recorded grids is expected to consist of two components: cosmic-ray background fluctuations and potential PBH bursts in the data.
@@
GRB 221009A is an extraordinarily bright burst and has been observed by LHAASO [41, 42]. In this work, we also detect GRB 221009A significantly using the blind scan method (details can be found in Supplemental Material [33]).
@@
As no significant PBH burst signal is observed in the -value distribution of the data, the upper limit on the local PBH burst rate density is determined at the 99 confidence level (CL). In our search, the nearby grids have approximately 80 overlap in both spatial and temporal searches. This overlap causes the grids to be nonindependent, altering the statistical behavior of each -value bin in Fig. 1.
@@
Hence, we select three low-background-level -value regions as a single -value bin for the three search durations to determine the upper limits of the local PBH burst density. The chosen -value regions are for s, for s, and for s, respectively. These ranges are optimized for the best sensitivity of PBH signals based on expected background counts and the simulated signal PDF. The details are described in Supplemental Material [33].
The probability of the observed grid counts in the three chosen -value regions is calculated by summing over the probabilities of all possible signal and background count combinations. Therefore, the likelihood with a local burst rate density is given by
where is the observed count of grids in the given -value regions for the data. represents the simulated signal probability for signal grid counts under the local burst rate density of , and denotes the probability for the remaining background counts.
To determine the 99 CL upper limits of the local burst rate density, we perform a likelihood scan under various hypotheses for the local burst rate density. After incorporating systematic uncertainty, the 99 CL upper limits on the local PBH burst rate density are listed in Table 1 and shown in Fig. 2. Compared to previous constraints from other telescopes and observatories, LHAASO demonstrates the highest sensitivity to timescales ranging from several seconds to several hundred seconds.
@@ 99 CL Upper Limit (pc-3 yr-1)
For background estimation, the detector acceptance is estimated using a three-hour integration time, and the background in each grid is then calculated.
To assess the uncertainty in the background histograms, the integration time is varied from three to five hours. Furthermore, we use the -value range of as the fiducial region to verify the consistency between the data and the background simulation. The systematic uncertainty from the background estimation is estimated to be 5 across the three burst timescales.
In the PBH burst simulation, the atmospheric model in the MC simulation, along with variations in atmospheric conditions during operation, affects the detection efficiency, which is not properly modeled in the simulation. This uncertainty is estimated to be as large as 8 on the effective area [43]. Additionally, uncertainties in the PSF influence the number of signals within the given grids, thus impacting the -value distribution of PBH signals. The uncertainty in the simulated PSF is estimated to be by comparing the observed event profiles among Crab, Mrk 421, and Mrk 501 [43]. To estimate the systematic uncertainty due to the effective area and PSF simulation, we resimulate the PBH burst signals by scaling the effective area by 92 and enlarging the PSF by . The total systematic uncertainty is accounted for by varying the background level within its corresponding uncertainty and using the resimulated signal PDFs to repeat the upper limit measurement. The larger upper limit, compared to the nominal one, is adopted as the final result.